### v3

@@ II System Architecture
then the perturbation per step satisfies $|r^{\prime}_{t}-r_{t}|\leq\lambda$. Over the discounted horizon, and for a fixed policy $\pi$, the cumulative shift is bounded by
$|V^{\prime\pi}(s)-V^{\pi}(s)|\leq\frac{\lambda}{1-\gamma}.$ (6)
@@
$|A^{\prime\pi}(s,a)-A^{\pi}(s,a)|\leq\frac{2\lambda}{1-\gamma}.$ (7)
@@
Beyond stability, these evaluations instantiate two fundamental game-theoretic principles. The first is belief revision, which parallels Bayesian updating. Let $b_{t}$ denote the LLM’s belief score of cooperative success given prompt $\phi(s_{t})$. We model this as
$b_{t+1}=\frac{P(\phi(s_{t+1})\mid\mathrm{coop})\,b_{t}}{P(\phi(s_{t+1})\mid\mathrm{coop})\,b_{t}+P(\phi(s_{t+1})\mid\lnot\mathrm{coop})\,(1-b_{t})}.$ (8)
so that the LLM revises its cooperative judgment in response to evolving prompt content such as pending orders, role distribution, or elapsed time. This allows the agent to update expectations mid-task without retraining, while still remaining deterministic.
@@
$s_{t}=\begin{cases}+1&\text{if }a_{t}\in\mathcal{A}_{\text{coop}},\\ 0&\text{otherwise}.\end{cases}$ (9)
where $\mathcal{A}_{\text{coop}}$ is the set of actions contributing meaningfully to shared objectives. This ensures agents receive positive shaping signals only when their behavior advances cooperative goals, thereby discouraging selfish or unstable strategies and promoting long-term alignment.
Finally, to guarantee that shaping never dominates the task reward, $\lambda$ must satisfy
$\lambda\leq\min\!\left\{(1-\gamma)\,\delta_{V},\;\tfrac{1-\gamma}{2}\,\varepsilon_{\text{clip}}\,A_{\text{scale}}\right\}.$ (10)
@@
To operationalize these principles, we embed structured symbolic evaluation into the reinforcement learning loop through the LLM. A prompt generation module continuously encodes task-relevant context, such as the number of pending orders, the current timestep, and relative agent positions. This structured description is passed to a frozen LLM, which produces a binary cooperative judgment based on its logits for tokens such as “good” and “bad.” Formally, if the logit for “good” exceeds that for “bad,” a scalar bonus $\lambda$ is applied as in Eq. (1), yielding a shaped reward $r^{\prime}_{t}=r_{t}+\lambda$ (Eq. LABEL:eq:shaped_reward_simple). Otherwise, the base reward $r_{t}$ is retained. This ensures that feedback reflects the overall quality of collaboration rather than isolated action outcomes.
The shaped reward $r^{\prime}_{t}$ is then used to compute temporal-difference errors $\delta_{t}$ (Eq. 3) and advantages $\hat{A}_{t}$ via generalized advantage estimation (Eq. 4), allowing cooperative feedback to influence long-horizon credit assignment in PPO. By integrating symbolic evaluations in this way, the framework guides agents toward strategic behaviors that emphasize task flow, anticipation of partner actions, and recovery from errors.
To ensure stability, we bound the effect of shaping at every stage. Equation 5 guarantees that per-step perturbations remain within $|r^{\prime}_{t}-r_{t}|\leq\lambda$, while Eqs. 6 and 7 show that the cumulative shifts in value and advantage estimates are capped. These guarantees prevent destabilization during PPO training even under noisy or partial observations.
Beyond stability, the LLM feedback instantiates two game-theoretic principles. First, belief revision (Eq. 8) parallels Bayesian updating: the LLM revises its cooperative judgment deterministically as the prompt evolves with state changes such as order backlog or elapsed time. Second, strategic consistency (Eq. 9) enforces that shaping signals are positive only when actions belong to the cooperative set $\mathcal{A}_{\text{coop}}$, thereby discouraging selfish or unstable strategies. Finally, Eq. 10 constrains $\lambda$ so that shaping acts only as a regularizer, never overwhelming the task reward.
@@
As a result, agents dynamically adapt their policies mid-episode using high-level, interpretable feedback. This architecture bridges low-level reinforcement learning with symbolic reasoning, enabling robust coordination, adaptive role assignment, and resilience to uncertainty, including observation noise, partial observability, and real-time variability.
By linking symbolic judgment to policy adaptation, the framework naturally enforces game-theoretic alignment over time. Both PPO agents respond independently to the same symbolic signal and adapt their strategies to maximize cooperative feedback. As training progresses, the agents converge toward a stable configuration in which each agent’s role complements the other. This convergence satisfies the key condition of a Nash equilibrium between two agents: once both agents reach their equilibrium strategies, neither can improve their outcome by deviating unilaterally. Importantly, the framework does not solve for equilibrium analytically. Instead, equilibrium-like behavior emerges through iterative feedback and reward shaping, without any need for direct coordination or access to the co-agent’s internal state.
In summary, we have developed a new real-time shaping framework that integrates symbolic evaluation from frozen LLMs with policy optimization, enabling agents to adapt strategies dynamically in cooperative multi-agent settings. By fusing high-level symbolic reasoning with low-level reinforcement learning, our framework allows agents to revise behavior in real time without the need for retraining, message-passing, or centralized control. This approach offers several key advantages: (1) it provides interpretable, prompt-driven feedback that shapes coordination decisions at each step; (2) it enables agents to handle uncertainty, teammate variability, and perceptual noise through context-aware shaping; and (3) it facilitates convergence to equilibrium strategies via reward modulation rather than direct policy manipulation. The architecture grounds strategic behavior in game-theoretic principles such as belief revision and role alignment, supporting robust adaptation in dynamic environments. These properties make the system broadly applicable to settings where policies must remain flexible, resilient, and efficient under changing cooperative demands. We now evaluate its performance through a series of controlled simulations.
Figure 2: Custom Overcooked-AI “Burger Kitchen” scenario used in our experiments. Agent A (blue) retrieves raw patties from the fridge and cooks them on the Stove, while Agent B (green) assembles cooked patties with buns and condiments at the Burger Station and delivers the finished burger through the Delivery Window. Dashed arrows indicate task flow between stations. This scenario imposes strict temporal dependencies and spatial constraints that require real-time role specialization, dynamic coordination, and mid-episode strategy adjustment,exactly the challenges our PPO+LLM shaping framework is designed to address.

### v4

@@ II System Architecture
where $c_{t}$ denotes a bounded symbolic shaping signal produced by the frozen LLM and is distinct from the environment state $s_{t}$, then the per-step perturbation satisfies $|r^{\prime}_{t}-r_{t}|\leq\lambda_{s}$. Over the discounted horizon, and for a fixed policy $\pi$, the cumulative shift in value estimates is bounded by
$\big|V^{\prime\pi}(s)-V^{\pi}(s)\big|\leq\frac{\lambda_{s}}{1-\gamma}.$ (6)
@@
$\big|A^{\prime\pi}(s,a)-A^{\pi}(s,a)\big|\leq\frac{2\lambda_{s}}{1-\gamma}.$ (7)
@@
Beyond stability, these evaluations instantiate two fundamental game-theoretic principles. The first is belief revision, which parallels Bayesian updating.
Recall that the LLM produces a discrete symbolic judgment (e.g., “good” vs. “bad”) reflecting whether the agents’ joint behavior appears cooperative. We model this judgment using a latent binary variable
$z\in\{\text{cooperative},\;\text{non-cooperative}\}.$
Let $b_{t}\in[0,1]$ denote the LLM’s belief at time $t$ that the agents are cooperating, conditioned on the textual prompt $\phi(s_{t})$ extracted from the environment state. This belief is updated using a Bayesian revision rule:
$b_{t+1}=\frac{P(\phi(s_{t+1})\mid z)\,b_{t}}{P(\phi(s_{t+1})\mid z)\,b_{t}+P(\phi(s_{t+1})\mid\neg z)\,(1-b_{t})}.$ (8)
Here, $z$ denotes the latent event that the agents’ joint behavior is cooperative, while $\neg z$ denotes non-cooperative behavior. Operationally, these cases correspond to the LLM assigning greater logit mass to positive (cooperative) or negative (non-cooperative) symbolic evaluations, respectively. Thus, the sign of the LLM’s internal evaluation implicitly determines whether the observed prompt evidence supports cooperation or non-cooperation.
The quantity $b_{t+1}$ represents the LLM’s posterior belief at the next timestep that the agents are cooperating, updated in response to the newly observed prompt derived from $s_{t+1}$. This revision captures how symbolic feedback—such as stalled orders, role conflicts, or inefficient task allocation—modulates cooperative belief over time. Because the LLM is frozen and evaluated deterministically, $b_{t}$ evolves predictably without requiring any retraining of the underlying policy.
@@
$c_{t}=\begin{cases}+1&\text{if }a_{t}\in\mathcal{A}_{\text{coop}},\\ 0&\text{otherwise},\end{cases}\qquad c_{t}\in[-1,1].$ (9)
Here, $\mathcal{A}_{\text{coop}}$ denotes the subset of actions that the frozen LLM evaluates as contributing to cooperative task flow in state $s_{t}$. In practice, an action is included in $\mathcal{A}_{\text{coop}}$ when the LLM assigns higher logit mass to the token "good" than to "bad" for the prompt describing $(s_{t},a_{t})$. These are typically actions that progress the shared objective, maintain role specialization, prevent collisions, or advance pending orders. By restricting the shaping signal to this cooperative set, the framework rewards only behavior that meaningfully supports joint success, discourages unstable or selfish deviations, and strengthens long-horizon coordination.
@@
Putting these components together, PPO+LLM embeds structured symbolic evaluation directly into the reinforcement learning loop. At each timestep, a prompt generation module encodes task-relevant context (e.g., pending orders, timestep, and relative agent positions) into a fixed text template that is scored by a frozen LLM using the logits of tokens such as “good” and “bad.” If the logit for “good” exceeds that for “bad,” a shaping bonus $\lambda_{s}$ is applied as in Eq. (1), yielding the shaped reward (Lemma 1); otherwise, the base reward $r_{t}$ is retained. The shaped reward $r^{\prime}_{t}$ then propagates through the standard PPO pipeline via the TD error and GAE (Eqs. (3)–(4)), allowing symbolic feedback to influence long-horizon credit assignment.
Stability is preserved by construction. For instance Eq. (5) bounds the per-step perturbation $|r^{\prime}_{t}-r_{t}|\leq\lambda_{s}$, while Eqs. (6) and (7) bound the induced shifts in value and advantage estimates. Finally, the clipping-aware constraint in Eq. (2) restricts $\lambda_{s}$ so that shaping acts only as a small regularizer and cannot dominate the task reward. Beyond these guarantees, the same mechanism instantiates two game-theoretic principles: belief revision via Eq. (8) and strategic consistency via Eq. (9).
@@
As a result, agents dynamically adapt their policies mid-episode using high-level, interpretable feedback. This architecture bridges low-level reinforcement learning with symbolic reasoning, enabling robust coordination, adaptive role assignment, and resilience to uncertainty such as observation noise, partial observability, and real-time variability. In practice, this symbolic feedback is applied symmetrically across agents while preserving independent learning dynamics.
Both agents receive the same symbolic feedback signal, but they process it independently during training. The signal is generated externally by the shaping module (not by either agent) and is added to each agent’s observation in the same way that standard shaped rewards are applied in multi-agent RL. Because the agents do not access one another’s internal states or intentions, they adapt their policies separately. Over training, these independent updates lead the agents to settle into complementary roles. Once this stable configuration emerges, no agent can improve its outcome by deviating alone, reflecting an approximate Nash equilibrium induced by the shaping dynamics.
To ensure architectural fairness, all agents share an identical PPO policy backbone with fixed representational capacity; consequently, any observed differences in coordination, robustness, or equilibrium behavior arise from system-level mechanisms such as symbolic shaping and environmental perturbations rather than from differences in model size or expressivity
II-A Approximate Nash Equilibrium Probe
To quantify approximate equilibrium behavior, we evaluate the stability of learned joint policies against unilateral deviations via a best-response Nash-gap probe. We perform this probe on the Overcooked-AI layout asymmetric_advantages, which consists of two separated kitchen areas with asymmetric access (different distances) to key task objects. This asymmetry induces natural specialization and complementary roles, making it a useful testbed for equilibrium-style stability under cooperative learning.
II-A1 Cooperative Markov game setting
We model Overcooked-AI on this layout as a two-agent cooperative Markov game with finite horizon $H=400$. At each timestep $t\in\{0,\dots,H-1\}$, the environment is in state $s_{t}\in\mathcal{S}$, agents select primitive actions $a_{t}^{1},a_{t}^{2}\in\mathcal{A}$, and the environment transitions according to
$s_{t+1}\sim P(\cdot\mid s_{t},a_{t}^{1},a_{t}^{2}),$ (10)
emitting a shared team reward
$r_{t}=r(s_{t},a_{t}^{1},a_{t}^{2}).$ (11)
Both agents share the same primitive action set, so $\mathcal{A}_{1}=\mathcal{A}_{2}=\mathcal{A}$. A joint policy $\pi_{\theta}$ maps states to distributions over joint actions,
$\pi_{\theta}(a^{1},a^{2}\mid s)\in\Delta(\mathcal{A}_{1}\times\mathcal{A}_{2}),$ (12)
and in our implementation is represented by a single PPO policy that outputs a two-dimensional discrete action vector corresponding to both agents’ primitive actions. The episodic return of a joint policy is
$V(\pi_{\theta})=\mathbb{E}\Bigg[\sum_{t=0}^{H-1}r_{t}\,\Bigg|\,a_{t}^{1},a_{t}^{2}\sim\pi_{\theta}(\cdot\mid s_{t})\Bigg].$ (13)
Our analysis assumes a family of trained joint PPO policies indexed by baseline label $b\in\mathcal{B}$, environment regime $e\in\mathcal{E}$, and random seed $s\in\mathcal{S}_{\mathrm{train}}$, where
$\mathcal{B}$ $=\{\text{Baseline},\,\text{PPO+LLM},\,\text{CC\_PPO},$
$\quad\text{SP\_PPO},\,\text{HARL},\,\text{PBT\_PPO}\},$ (14)
$\mathcal{E}$ $=\{\text{No Noise},\,\text{Noise},\,\text{Delay},\,\text{Combo}\},$ (15)
$\mathcal{S}_{\mathrm{train}}$ $=\{1001,2002,3003,4004,5005\}.$ (16)
II-A2 Unilateral best-response environment
To probe stability against unilateral deviations, we construct a single-agent best-response environment that holds one agent fixed to the learned joint policy $\pi_{b,e,s}$ while allowing the other agent to deviate. Concretely, the learning agent observes the same flattened Overcooked feature representation as in training, but selects a single primitive action index $a_{t}^{\mathrm{BR}}\in\{0,\dots,|\mathcal{A}|-1\}$ for the controlled player.
At each timestep, the fixed opponent produces its action via deterministic inference from the frozen joint policy. The environment then forms the joint action by combining (i) the learning agent’s chosen action for the controlled index and (ii) the frozen opponent’s action for the other index, and applies the resulting joint action to the underlying Overcooked simulator. This induces a single-agent MDP $M_{\mathrm{BR}}(\pi_{b,e,s})$ in which the dynamics and team reward are inherited from Overcooked-AI, one player is controlled by the learning agent, and the other follows the fixed policy.
II-A3 Nash gap computation
Let $V_{\mathrm{self}}^{b,e,s}$ denote the self-play value of the joint policy $\pi_{b,e,s}$, and let $V_{\mathrm{BR}}^{b,e,s}$ denote the value achieved when one agent is replaced by an approximate best response trained against the frozen opponent induced by $\pi_{b,e,s}$. We define the Nash gap as
$\Delta^{b,e,s}=V_{\mathrm{BR}}^{b,e,s}-V_{\mathrm{self}}^{b,e,s}.$ (17)
The interpretation is:
  • If $\Delta^{b,e,s}>0$, then a unilateral deviation increases return, indicating exploitable behavior.
  • If $\Delta^{b,e,s}<0$, then the deviation underperforms self-play, suggesting local stability with respect to the deviation class.
  • If $\Delta^{b,e,s}\approx 0$, then the policy behaves as an approximate Nash equilibrium relative to the best-response procedure used.
In practice, for each $(b,e,s)$ we (i) evaluate self-play to estimate $V_{\mathrm{self}}^{b,e,s}$, (ii) train an approximate best response against the frozen opponent induced by $\pi_{b,e,s}$ to estimate $V_{\mathrm{BR}}^{b,e,s}$, and (iii) report $\Delta^{b,e,s}$. Implementation details (wrapper construction, self-play evaluation protocol, and best-response training procedure) are provided in the appendix.